can you please help to solve the following for x
 
solve sin^-1(6x)+sin^-1(6sqrt(3)x)=-pi/2

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Devindar
 
pi/2 + sin^-1(6 sqrt(3) x) = - sin^-1(6x)
 
Taking sin both the sides
 
Cos (sin^-1(6sqrt(3) x)) = -6x
 
On solving further
x = 1/12 and  -1/12
 
But x = 1/12 doesn't satisy hence
x = -1/12 is tye Correct answer.
 
Regards
Arun (askIITians forum expert)